find the least common denominator
3/4,2/5

2/9,3/18

1/16,9/20

14/54,31/81


please solve these four question
soon
i really need help in them

Answer by stanbon(75887)

find the least common denominator
3/4,2/5
--
lcd = 4*5 = 20
==================
2/9,3/18
Since 18 = 2*9
lcd = 18
==================
1/16,9/20
16 = 2^4
20 = 2^2*5
lcd = 2^4*5 = 80
==================
14/54,31/81
54 = 2*3^3
81 = 3^4
lcd = 2*3^4 = 162
=========================
Procedure for find the lcd of a set of numbers:
1st: Write all the denominators in prime-factor form
2nd: The lcd is the product of each of the prime
factors to their highest power.
---
Example:
14/54,31/81
54 = 2*3^3
81 = 3^4
--
2 is to the 1st power
3 is to the 4th power
-------------------------
lcd = 2*3^4 = 162
Cheers